# Neuromodulation and homeostasis: complementary mechanisms for robust neural function

### v1

@@ ABSTRACT
Neurons depend on two interdependent mechanisms—homeostasis and neuromodulation—to maintain robust and adaptable functionality. Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals. Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning. This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function. Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while maintaining intracellular calcium levels. Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades. The interaction between these controllers critically depends on the existence of an intersection in conductance space, representing a balance between target calcium levels and neuromodulated firing patterns. Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades. We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators in crustaceans. These findings suggest that targeting neuromodulation pathways—rather than ion channels directly—may offer safer pharmacological strategies to manage neuronal dysfunctions. This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
SIGNIFICANCE STATEMENT
Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals. These mechanisms often interact, but their improper coordination can lead to dysfunction. This study uses computational models to show how controlled neuromodulation - mimicking biological feedback - can be harmonized with homeostasis to ensure robust neuronal function. By improving neuronal degeneracy, the system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks. These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
@@ INTRODUCTION
Brain activity is continuously shaped by neuromodulators and neuropeptides, including dopamine, serotonin, and histamine (Bargmann and Marder,, 2013; Marder et al.,, 2014). Neuromodulators dynamically influence single-neuron activity, input-output properties, and synaptic connection strength and dynamics, enabling neuronal networks to adapt to changing needs, contexts, and environments (Marder and Calabrese,, 1996; Marder and Bucher,, 2001; McCormick et al.,, 2020). To achieve this modulation, neuromodulators dynamically reshape the density, dynamics, and kinetics of many single cell transmembrane proteins, making whole brain functional signaling strongly dependent on the robustness and reliability of neuromodulation mechanisms at the molecular and cellular levels (Nadim and Bucher,, 2014; Hooper and Marder,, 1984). While experimental studies underscore the ubiquity of neuromodulation across nervous systems, the fundamental principles underlying this mechanism remain incompletely understood. As such, studying neuromodulation is a very arduous task, both experimentally and computationally, which explains our limited understanding of its underlying mechanisms.
@@
Neuromodulation coexists with cellular homeostasis, a mechanism that gradually adjusts neuronal membrane properties to maintain a target activity level (Desai,, 2003; Marder and Goaillard,, 2006; O’Leary et al.,, 2013). Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals (Marder,, 2011). Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks (O’Leary and Wyllie,, 2011), the robustness and excitability of neurons must remain undisturbed (Marder and Prinz,, 2002). By monitoring only intracellular calcium levels, the neuron orchestrates the regulation of all its ion channel conductances through homeostatic self-tuning rules (O’Leary et al.,, 2010; Pratt and Aizenman,, 2007). This raises the question of how two different mechanisms acting on the same targets (ion channels), each potentially having divergent goals, can co-exist and complement each other harmoniously (Marder et al.,, 2014; Marder and Prinz,, 2002).
We propose to study this interaction between neuromodulation and homeostasis computationally, using conductance-based models. Conductance-based models provide an electrical representation of the neuronal membrane, where the membrane is modeled as an RC circuit, while the ion channels are modeled as non-linear conductances (Izhikevich,, 2007). These models are the closest equivalent of biological neurons, thanks to their ability to model any ion channel from experimental data following a methodology first proposed by Hodgkin and Huxley (Hodgkin and Huxley,, 1952). There exists a myriad of different conductance-based models that model various types of neuronal cells (Liu et al.,, 1998; Qian et al.,, 2014; Plant,, 1981; Wechselberger et al.,, 2006; Drion et al.,, 2011; Katori et al.,, 2010; Megwa et al.,, 2023).
On the one hand, a homeostatic controller has been developed by O’Leary et al., (2014) and can be applied on top of any conductance-based model. In this controller, the neuron can adjust all ionic conductances up or down depending on its average intracellular calcium level. If the calcium level is lower than a reference value, ionic conductances are increased following homeostatic tuning rules and vice versa. Such rules force the tuning of ionic conductances along a single axis, that is, by preserving conductance ratios or correlations. On the other hand, Fyon et al., (2023) has developed a reliable neuromodulation controller and can also be applied to any conductance-based model. This controller can robustly tune the external activity of neurons, which means that neurons with various sets of ion channel properties that lead to similar firing activities (a property called degeneracy (Drion et al., 2015b, )) can be neuromodulated using the same control action. In this study, we first examine the interaction between homeostasis and sharp neuromodulation, emphasizing that improperly combining the homeostatic controller with neuromodulation may result in unreliable outcomes. Subsequently, we expand our analysis to explore the interaction between the homeostatic controller and a neuromodulation controller, both at the single-cell level and in rhythmic networks. This approach enables the development of a combined controller that integrates the strengths of both individual controllers, facilitating robust modulation and preservation of neuronal function.
@@ RESULTS
Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation
This study employs a stomatogastric ganglion (STG) neuron conductance model (Liu et al.,, 1998) to explore the interaction between homeostasis and neuromodulation. These neurons contribute to rhythmic circuits that generate signals triggering stomach muscle contractions in crustaceans. This model was chosen for its ability to produce diverse firing patterns—ranging from tonic spiking to bursting, including single-spike bursting—and its rich combination of ion channels enabling degeneracy. The model includes classical fast sodium and delayed rectified potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels as well as hyperpolarization-activated H-channels. We first study the interaction between state-of-the-art model homeostasis (O’Leary et al.,, 2014) and sharp neuromodulation (Fyon et al.,, 2024). We define sharp neuromodulation as a direct change in a subset of ion channel maximal conductances, which results in functional modulation. In the context of this work, we will focus on a robust transition from tonic spiking to bursting in degenerate neurons, i.e. neurons exhibiting similar firing activities from different underlying channel conductances.
First, we sharply neuromodulated a set of degenerate neurons from tonic spiking to bursting by directly modifying the maximal conductances of A-type potassium $\bar{g}_{A}$ and slow calcium $\bar{g}_{CaS}$, then allowed homeostatic compensation to take its course. Three example traces with a time evolution of maximal conductances are provided in Fig. 1, left, whereas Fig. 1, right show trajectories in the $\bar{g}_{\mathrm{CaS}}$-$\bar{g}_{\mathrm{A}}$ plane. As can be seen in the figure, although the initial tonic firing and the sharply neuromodulated bursting activities are very similar in all three neurons, homeostatic compensation has very variable and unreliable effects. Depending on the initial conductance state, neuromodulation shifts the neuron to a new state where homeostatic compensation can lead to pathological (Fig. 1A), functionally preserved (Fig. 1B), or unphysiological (Fig. 1C) behaviors. This interesting result highlights an important property of neuronal degeneracy: although distinct channel subsets can produce similar neuronal behaviors, neurons may respond differently to perturbations, such as sharp neuromodulation.
To better understand the reasons behind this unreliability, both mechanisms must be discussed in detail. In the homeostatic model O’Leary et al., (2014), the cell senses its own intracellular calcium level and averages it over time. If the mean intracellular calcium falls below a predefined target value, RNA transcription and ion channel protein translation are triggered to increase all ionic conductance values, and vice versa. Specifically, increasing calcium ion channel conductances leads to more calcium ions entering the cell. This implies that homeostasis induces changes in all conductance values (Fig. 1 left), which leads to a linear motion in the conductance plane (Fig. 1 right). This phenomenon happens on a relatively slow timescale, which can last up to days (O’Leary et al.,, 2010). Notably, the RNA transcription time constants, denoted by $\tau_{m}$, vary across different ion channels and cell types. Steady-state mathematical analysis has shown that these differing time constants tune correlations in ion channel conductances. For two ion channels $i$ and $j$, steady-state analysis yields $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$. This indicates that homeostatic tuning rules maintain ratios of ionic conductances and their correlations while adjusting their values in response to intracellular calcium levels. In the ionic conductance space, these rules allow neurons to navigate along a one-dimensional line, where the slope of this line in two dimensions is determined by the ratios of mRNA transcription time constants. This phenomenon, known as homogeneous scaling, involves proportionally increasing or decreasing all ionic conductances by the same multiplicative factor.
On the other hand, sharp neuromodulation is implemented by transiently applying the algorithm of Fyon et al., (2024). This neuromodulation also produces changes in conductance values (Fig. 1 left), resulting in motion within the conductance plane in a direction different from homogeneous scaling (Fig. 1 right) to reach a new firing activity. However, neuromodulation only affects a subset of ion channels, not all (Marder,, 2012). Furthermore, the timescale of neuromodulation is much faster than that of homeostasis, lasting for a maximum of a few minutes (Marder et al.,, 2014). In addition, neuromodulation tunes correlations between neuromodulated ion channels, rather than preserving them (Fyon et al.,, 2024).
Coupling homeostasis and neuromodulation in this way is problematic because homeostasis preserves correlations and regulates calcium, whereas neuromodulation adjusts correlations to modulate neuronal function. In Fig. 1, sharp neuromodulation is applied to three different STG models with distinct initial ionic conductances. Initially, all neurons exhibited tonic firing due to homeostatic regulation, with RNA transcription time constants selected to produce tonic spiking neurons. Neuromodulation was then applied instantaneously, leading to a transition from tonic spiking to bursting in all cases (Fig. 1). These transitions drastically increase intracellular calcium concentration. From the perspective of homeostatic tuning rules, this constitutes a disturbance, as it alters the conductance ratios between $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$. Although this disturbance is necessary for adaptation and functional modulation, the outcome of homeostasis in response to this biological disturbance is unpredictable. Indeed, as the homeostatic controller was initially designed to produce tonic spiking neurons, it tends to preserve the predefined correlations that produce such behavior. However, neuromodulation changes the firing activity, leading to a mismatch between actual conductance values and the encrypted conductance ratios.
In Fig. 1A, homeostatic compensation results in pathological behavior, characterized by irregular bursting with a hyperpolarized plateau, failing to achieve either bursting or tonic spiking. In Fig. 1B, the modulated behavior (bursting) is maintained after homeostatic compensation. In Fig. 1C, homeostatic compensation leads to unphysiological values of $\bar{g}_{\mathrm{A}}$, tending towards negative values. Even when $\bar{g}_{\mathrm{A}}$ is constrained to be non-negative, the resulting electrical behavior is pathological, failing to exhibit either tonic spiking or bursting. This toy experiment highlights that homeostasis and sharp neuromodulation can interfere pathologically.
Such behavior was predicted in O’Leary et al., (2014), where the authors demonstrate that their homeostasis model can lead to various outcomes in response to disturbances, such as channel deletions. Initially, the deletion itself may directly affect neuronal function, either preserving it or causing its loss. Subsequently, homeostatic compensation can either preserve or restore function, but it may also lead to pathological behaviors, such as functional deletion. This result suggests that any disturbance applied to the conductance ratios—even physiological ones such as neuromodulation—can lead to problematic behaviors.
Figure 1: Current model of homeostasis can lead to pathological excitability in the presence of sharp neuromodulation. A. Time evolution of all conductances in the STG model, shown on a logarithmic scale, when homeostasis is applied with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ within the red region) (left). Corresponding trajectory in the modulated conductance space (right). The white circle represents the initial condition (tonic spiking), the red diamond marks the neuromodulated state (bursting) immediately after the sharp modulation, and the green triangle denotes the final state after homeostasis has converged. In this case, homeostatic compensation is pathological, resulting in irregular bursting. B. Same as in panel A, but with different degenerate initial conditions. Here, homeostatic compensation preserves the neuromodulated function. C. Same as in panel A, with another set of degenerate initial conditions. In this case, homeostatic compensation results in negative values of $\bar{g}_{\mathrm{A}}$ (saturated at 0), leading to pathological behavior characterized by one-spike bursting.
Controlled neuromodulation and homeostasis naturally lead to robust neuronal function
In biological systems, homeostasis and neuromodulation collaborate to ensure robust and modular neural function. These mechanisms operate on different timescales: neuromodulation acts within seconds or minutes (Werner and Mitterauer,, 2014), while homeostatic compensation may take hours or even days (Marder and Prinz,, 2002). However, in computational models, the goals of neuromodulation and homeostatic controllers often seem discordant. Neuromodulation typically adjusts conductance ratios, while homeostasis seeks to maintain these ratios and guide the neuron through homogeneous scaling. Consequently, applying sharp neuromodulation to any homeostatic conductance-based model can lead to instability, as neuromodulation disrupts the references and control actions of the homeostasis controller (Fig. 2A).
To model the actions of both mechanisms within a single cell, we include a more advanced neuromodulation model, referred to as the neuromodulation controller, as developed in Fyon et al., (2023). This controller can be coupled with any conductance-based model, allowing modulation of its neuronal function, such as its firing pattern, to mimic biological neuromodulation. We call this type of neuromodulation controlled neuromodulation, in contrast to the sharp neuromodulation discussed above. The neuromodulation controller adjusts target channel conductances, providing continuous and consistent modulation rather than instantaneous, rigid changes.
Fig. 2B illustrates the results of transitions from tonic spiking to bursting when applied to a degenerate population of STG neuron models. Initially, all neurons exhibit tonic firing mode, and homeostasis adjusts the conductances to achieve a predefined calcium level along a first tonic spiking homogeneous scaling line. Midway through the simulation, controlled neuromodulation is applied to switch all neurons to bursting mode, which now occurs through a controlled adjustment of $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$. Consequently, intracellular calcium levels rise sharply, and conductance ratios involving one or both neuromodulated ionic conductances are adjusted. Finally, with the reshaped RNA time constants, the homeostasis controller adjusts all conductances to restore an appropriate calcium level, while controlled neuromodulation preserves neuromodulated conductance ratios along a second bursting homogeneous scaling line. Ultimately, the neuronal population reaches a new state where the neuromodulated function is preserved and intracellular calcium levels are back to their target value.
Simulations were also performed on a midbrain dopaminergic neuron model (adapted from Qian et al., (2014)) to demonstrate the generality of our results (Fig. 2C). As with the previous simulations, a transition from tonic spiking (pacemaking) to strong bursting was induced midway through by modulating the N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) using controlled neuromodulation. Homeostatic compensation was continuously applied to maintain proper calcium levels, resulting in a reliable and robust outcome despite initial degeneracy within the population.
@@
Figure 2: Controlled neuromodulation and homeostasis naturally lead to robust neuronal function. A. Schematic representation of neuromodulation (left) and homeostasis (right) cascades. Neuromodulators bind to G-protein receptors, triggering complex signaling pathways that selectively affect subsets of ion channels. In contrast, homeostasis senses intracellular calcium levels and regulates them by adjusting all conductances uniformly. B. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (left). The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation. C. Same as panel B, but for a DA model where controlled neuromodulation affects $\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$.
@@
Why do controlled neuromodulation and homeostasis lead to robust neural function?
@@
Fig. 3A shows the trajectories of the simulations from Fig. 2B in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation. Initially, the modulated conductances of every neuron in the population grow along the homogeneous scaling direction in tonic spiking mode, as the calcium levels are below the threshold value. This homogeneous scaling corresponds to a line intersecting the origin, with variability due to the degeneracy in conductance ratios within the population. During neuromodulation, all conductances are adjusted to tune their ratios and achieve bursting (Fig. 3B). Importantly, all neuromodulatory actions occur along the same direction, but with varying magnitudes depending on the state of the neuron (Fyon et al.,, 2024). As a result, calcium levels increase drastically for each neuron in the population. Once the neuromodulatory action ends, the outcome depends on the type of neuromodulation applied (Fig. 3C).
In the case of sharp neuromodulation, during the homeostatic action in the bursting phase, the homeostatic controller acts alone and tends to maintain the same conductance ratios as in tonic spiking (Fig. 3C left). This leads to a reduction in calcium levels, but, as seen in Fig. 1, its effects are unreliable and undesirable. Specifically, for most neurons in the population, the homeostatic controller reduces $\bar{g}_{\mathrm{A}}$ towards negative values (saturated at 0 to avoid model instability) to bring $\bar{g}_{\mathrm{CaS}}$ into an appropriate range. This regulation causes the neuron to move away from the robust bursting behavior, resulting in heterogeneous changes in the activity of the population.
In the case of controlled neuromodulation, the homeostatic and neuromodulation controllers work together to maintain the new conductance ratios corresponding to bursting (Fig. 3C right). This results in a new homogeneous scaling following these updated conductance ratios in bursting, which translates into trajectories that move toward the origin of the axes, as would be expected in tonic spiking. Homogeneous scaling during bursting differs across the population due to initial degeneracy. The difference between the two homogeneous scalings (tonic spiking and bursting) is driven by the neuromodulatory action, ensuring that the neuromodulated behavior (bursting) is not lost.

### v2

# Controlled neuromodulation naturally interacts with homeostasis to produce robust and modulable neuronal function

@@ Abstract
Neurons rely on two interdependent mechanisms – homeostasis and neuromodulation – to maintain robust and adaptable functionality. Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals carried by neuromodulators. Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning. This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function. Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while homeostasis simultaneously maintains target intracellular calcium levels. Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades. The interaction between these controllers critically depends on the existence of an intersection in conductance space, representing a balance between target calcium levels and neuromodulated firing patterns. Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades. We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators rhythmic activity. This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
Author summary
Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals. These mechanisms often interact, but their improper coordination can lead to dysfunction. This study uses computational models to show how controlled neuromodulation – mimicking biological feedback – can be harmonized with controlled homeostasis to ensure robust neuronal function. The system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks even when faced with neuronal parameter variability. These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
@@ Introduction
Brain activity is continuously shaped by neuromodulators and neuropeptides, including dopamine, serotonin, and histamine [2, 30]. Neuromodulators dynamically influence single-neuron activity, input-output properties, and synaptic strength, enabling neuronal networks to adapt to changing needs, contexts, and environments [28, 27, 33]. This modulation occurs through the regulation of transmembrane proteins, such as ion channels and receptors, which alters neuronal excitability and synaptic dynamics [35, 20]. While experimental studies highlight the ubiquity of neuromodulation, the fundamental principles governing its effects remain incompletely understood, making computational approaches crucial for unraveling its mechanisms.
@@
Neuromodulation coexists with homeostatic plasticity, a process that gradually adjusts neuronal membrane properties to maintain a target activity level [8, 29, 37, 7, 50, 49, 51]. Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals [32]. Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks [39], the robustness and excitability of neurons must remain undisturbed [31]. Among homeostatic mechanisms, intrinsic homeostatic plasticity adjusts ion channel conductances in response to intracellular calcium levels [36, 41]. However, since both neuromodulation and homeostasis act on the same targets (ion channels) and may have competing objectives, understanding their interaction is essential [30, 31].
We propose to study the interaction between neuromodulation and homeostasis using conductance-based models. These models provide a biophysically accurate representation of neuronal dynamics by describing the electrical properties of the neuronal membrane. The membrane is modeled as an RC circuit, while ion channels are represented as nonlinear conductances that evolve according to experimentally derived activation and inactivation kinetics [22]. Conductance-based models remain the most faithful mathematical representations of biological neurons, as they allow for the direct incorporation of ion channel dynamics extracted from experimental data, following the methodology first introduced by Hodgkin and Huxley [19]. Over the years, numerous conductance-based models have been developed to describe a variety of neuronal cell types across different species and brain regions [25, 43, 40, 52, 12, 23, 34]. This study focuses on a conductance-based model of stomatogastric ganglion (STG) neurons [25]. STG neurons are key components of rhythmic circuits that generate motor patterns driving stomach muscle contractions in crustaceans. This model was chosen for its ability to reproduce diverse firing patterns – ranging from tonic spiking to bursting – and its rich set of ion channels, which enable degeneracy. Specifically, the model includes classical fast sodium and delayed rectifier potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels, as well as hyperpolarization-activated H-channels.
We model intrinsic homeostasis by using the homeostatic controller proposed by [38], which adjusts all ionic conductances up or down based on the neuron average intracellular calcium level. If the calcium level is lower than a target reference value, the controller increases ionic conductances according to homeostatic tuning rules, and conversely, decreases them if the calcium level is above the target. Importantly, these rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them. This self-tuning mechanism allows neurons to counteract perturbations while preserving stable activity.
Neuromodulation is modeled using two different approaches. On the one hand, we model neuromodulation as an acute change in target conductance values leading to a desired neuromodulated state in degenerate populations, which we call ”sharp neuromodulation”. On the other hand, we implement the neuromodulation controller developed by [14], which is designed to enable robust modulation of neuronal activity. We call this approach ”controlled neuromodulation”. One of its key properties is that it accounts for neuronal degeneracy, meaning that neurons with different ion channel compositions but similar activity patterns respond in a consistent manner to neuromodulation [15]. This feature is particularly important because biological neuronal populations exhibit significant variability in their underlying conductance parameters while maintaining functional stability. By leveraging degeneracy, the neuromodulation controller can apply the same control signal across a population of neurons with diverse conductance profiles, ensuring reliable and predictable neuromodulation.
We first investigate the interaction between intrinsic homeostatic plasticity and neuromodulation at the single-neuron level and show that combining intrinsic homeostasis with either sharp or controlled neuromodulation leads to vastly different outcomes. Secondly, we provide a mechanistic explanation for these different outcomes. We then extend our analysis to rhythmic neuronal networks, where neuromodulation plays a critical role in adjusting oscillatory patterns. By carefully integrating homeostatic and neuromodulatory control mechanisms, we aim to develop a unified framework that combines the strengths of both approaches. This integrated controller enables robust modulation of neuronal activity while ensuring long-term stability, shedding light on how neuromodulation and homeostasis coexist in biological systems.
@@ Results
Interference between homeostasis and sharp neuromodulation leads to pathological outcomes
We first examine the interplay between a state-of-the-art model of cellular homeostasis [38] and sharp neuromodulation [13]. Here, sharp neuromodulation is defined as a direct, externally driven modification of a subset of ion channel maximal conductances, resulting in a change in neuronal spiking behavior. We specifically focus on a robust transition from tonic spiking to bursting in degenerate neuronal populations, i.e. populations composed of neurons that exhibit similar firing behaviors despite differences in underlying conductance profiles.
To explore this interaction, we applied sharp neuromodulation to a population of degenerate neurons generated using the dataset from [13] (Fig 1). Sharp neuromodulation consisted of directly modifying the maximal conductances of A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) channels to transition neurons from tonic spiking to bursting. We then allowed homeostatic compensation to unfold. Since bursting increases overall neuronal activity, intracellular calcium levels rise, triggering a decrease in all conductances according to homeostatic tuning rules. Three representative conductance trajectories ($N=200$) are shown in Fig 1 (top), while Fig 1 (bottom) illustrates calcium dynamics across the population (note that to facilitate observation of both mechanisms within a single trace, the homeostatic controller was accelerated compared to its original implementation). Despite initially similar firing patterns and a similar outcome of sharp neuromodulation, homeostatic compensation produced highly variable and often unreliable outcomes in terms of modulated activity, while still regulating intracellular calcium. Depending on the initial conductance state, neuromodulation led to new states where homeostasis could either maintain functional activity, result in pathological behavior, or drive neurons into unphysiological regimes. This finding underscores a fundamental property of degeneracy: neurons with different underlying conductances can exhibit similar activity yet respond differently to perturbations, such as neuromodulation.
@@
Fig 1: Interference between homeostasis and sharp neuromodulation leads to pathological outcomes. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (top). The corresponding mean intracellular calcium concentration over time (bottom) shows the target value (red dash-dotted line) is maintained both before and after sharp neuromodulation.
To understand the source of this unreliability, it is necessary to examine both mechanisms in more detail. The homeostatic model [38] relies on intracellular calcium as a feedback signal. The neuron continuously averages its intracellular calcium level over time and adjusts conductances accordingly: if calcium falls below a target level (red dash-dotted line in Fig 1 bottom), ion channel transcription and translation increase, raising all conductance values, and vice versa. This phenomenon happens on a relatively slow timescale, which can last up to days [36]. In contrast, sharp neuromodulation, implemented using the method from [13], directly modifies a subset of ion channel conductances, which upsets the homeostasis setpoint, defined by the intracellular calcium target level. As homeostasis kicks in to recover its calcium set point, there are no guarantees that the neuromodulated function is maintained as ion channel conductances are modified, which often leads to pathological or unphysiological outcomes.
This type of interference was anticipated in [38], where the authors demonstrated that homeostatic regulation can produce diverse responses to perturbations such as ion channel deletions. Depending on the initial conditions, a deletion may either preserve function, disrupt it, or lead to pathological compensation. Our results extend this idea by showing that even physiological disturbances, such as neuromodulation, can induce similar unpredictability when coupled with homeostatic regulation.
@@
Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function
In biological systems, neuromodulation occurs through an intracellular cascade that indirectly links changes in neuromodulation concentration and changes in targeted conductance values using an intermediate signal such as e.g. cyclic amp. We have recently shown that this indirect action of neuromodulation on conductance values could be viewed as an intracellular controller (Fig 2A, top), which largely improves reliability of neuromodulatory action on highly degenerate populations [14]. We call this mechanism ”controlled neuromodulation” to contrast with the sharp neuromodulation described above. Unlike sharp neuromodulation, which imposes instantaneous conductance changes, the neuromodulation controller provides continuous and adaptive modulation, ensuring stable functional transitions.
Fig 2: Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function.
A. Schematic representation of neuromodulation (left) and homeostasis (right) cascades. Neuromodulators bind to G-protein receptors, triggering complex signaling pathways that selectively affect subsets of ion channels. In contrast, homeostasis senses intracellular calcium levels and regulates them by adjusting all conductances uniformly.
B. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$) for a degenerate population of 200 models (left). The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation.
Using the same protocol as for sharp neuromodulation, we now study the interaction between controlled neuromodulation and homeostasis. Fig 2B illustrates the transition from tonic spiking to bursting in a degenerate population of STG neuron models, identical to those in Fig 1, but when using controlled neuromodulation instead of sharp neuromodulation. Initially, all neurons exhibit tonic firing, with homeostasis tuning conductances along a tonic spiking homogeneous scaling line to maintain a target calcium level. Midway through the simulation, controlled neuromodulation is applied, gradually adjusting $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ to induce bursting. As in the case of sharp neuromodulation, this transition increases intracellular calcium levels, leading to an adaptive shift in conductance values. Homeostatic regulation then readjusts overall conductance magnitudes, but here the interaction between homeostasis and controller neuromodulation ensures that the newly established neuromodulated state is maintained. Ultimately, the population reaches a stable configuration where both the neuromodulated function and intracellular calcium homeostasis are fulfilled.
To demonstrate the generality of this approach, we applied the same framework to a midbrain dopaminergic neuron model adapted from [43] (see Supporting Information, S1 Appendix). As in the STG model, a transition from tonic spiking (pacemaking) to strong bursting was induced by modulating N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) through controlled neuromodulation. Homeostatic compensation continuously regulated calcium levels, ensuring a reliable and robust outcome despite initial conductance degeneracy within the neuronal population.
Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation
@@
To study the mechanisms underlying the constructive interaction between controlled neuromodulation and homeostasis, we plot the conductance trajectories of the simulations from Fig 1 in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation (Fig 3A). Initially, as calcium levels remain below the threshold, homeostasis increases the modulated conductances along a homogeneous scaling direction characteristic of tonic spiking. This direction corresponds to a line intersecting the origin, with variability due to conductance degeneracy within the population, and its slope is determined by the relative time constants of mRNA transcription in the homeostasis model, with steady-state analysis predicting $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$ [38]. When neuromodulation is applied, all conductances are adjusted to reshape their ratios and induce bursting (Fig 3B). Importantly, neuromodulatory actions follow the same direction across neurons but vary in magnitude depending on individual neuronal states [13]. This modulation leads to a sharp increase in intracellular calcium levels. Once neuromodulation ceases, the long-term outcome depends on the type of neuromodulation applied (Fig 3C).
Fig 3: Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation.